# Why Rectified Power Unit Networks Fail and How to Improve It: An Effective Theory Perspective

### v3

@@ 6 Conclusion
In this paper, we analyzed neural networks using Effective Field Theory to theoretically predict and experimentally verify why neural network architectures with RePU activation fail when the network depth increases. Furthermore, we proposed the MRePU activation function, which improves upon RePU, and estimated its effectiveness theoretically. We then empirically confirmed its feasibility by comparing the experimental results with the theoretical predictions. The proposed MRePU was able to facilitate learning even in deep networks without adopting structures like ResNet. Mathematically, we confirmed that MRePU retains the favorable properties of RePU, such as p-differentiability and universal approximation properties. This suggests that the beneficial characteristics proposed in previous studies on RePU can also be extended to MRePU. Moreover, when attempting to approximate a function with a specific differentiability class, MRePU can be used to introduce inductive bias into deep neural networks, thereby enabling effective learning.

### v4

@@ 6 Conclusion
In this paper, we analyzed neural networks using Effective Field Theory to theoretically predict and experimentally verify the failure of deep neural network architectures employing RePU activation. Furthermore, we proposed the Modified Rectified Power Unit (MRePU) activation function to address the limitations of RePU and provided a theoretical estimation of its effectiveness. We then empirically validated its feasibility by comparing experimental results with theoretical predictions.
The proposed MRePU demonstrated the ability to facilitate learning even in deep networks. Mathematically, we confirmed that MRePU retains the favorable properties of RePU, such as $p$-differentiability and universal approximation capabilities. This suggests that the beneficial characteristics identified in previous studies on RePU can be extended to MRePU. Moreover, MRePU introduces a specific inductive bias into deep neural networks, thereby enabling effective learning when approximating functions within a specific differentiability class.
We also verified that the derived criticality condition aligns well with empirical validation on the MNIST dataset, a standard benchmark in computer vision, proving the effectiveness of MRePU in terms of both the stability of training dynamics near the fixed point and test accuracy. Furthermore, we extended our experiments to the more challenging CIFAR-10 task using a 14-layer ResNet architecture. While RePU activation failed to train at this depth, MRePU successfully demonstrated the capability to train on such deep and complex datasets.